OpenStudy (anonymous):

I will award you! :) For which system of equations is (2, 2) a solution? A. –3x + 3y = 0 x + 6y = 10 B. –2x + 5y = –6 4x – 2y = 4 C. 5x – 2y = –6 3x – 4y = 2 D. 2x + 3y = 10 4x + 5y = 18

OpenStudy (anonymous):

for this problem all you have to do is plug in your coordinate values into the equations to find the solution. So you'd just put in 2 for x and 2 for y. But you have to make sure it works for BOTH equations.

OpenStudy (anonymous):

D matches the requirements. 2(2)+3(2)= 10 4(2)+5(2)=18

OpenStudy (anonymous):

The correct answer would be D because the values plugged in equal the correct values. \[2(2)+3(2)=10\] \[4(2)+5(2)=18\]

OpenStudy (anonymous):

Thank you both! :)